FIND DOMAIN of following function:
1) f(x)= 1/x+[x]
2) f(x)=√(sinx-1)
please provide with procedure.
1
Expert's answer
2012-03-20T12:11:47-0400
1) x is in denominator, so x≠0 and D = R\{0}.
2) sinx-1 is under the arithmetic square root, so sinx-1 >= 0 ==> sinx >=1 ==> x = π(1+2n) and D = {x = π(1/2+2n), nєZ}.